OpenStudy (anonymous):

Using the quadratic formula to solve 5x = 6x2 – 3, what are the values of x?

OpenStudy (shamim):

6x^2-5x-3=0 Right?

OpenStudy (shamim):

U know a stanrd quadrutic equation is ax^2+bx+c=0

OpenStudy (shamim):

Solution x=(-b+-root(b^2-4ca))/2a

OpenStudy (shamim):

In ur given equation a=6 b=-5 c=-3

OpenStudy (shamim):

x=?

OpenStudy (hockeychick23):

5x=6x^2-3 --> 5x-5x=6x^2-5x-3 6x^2-5x-3 formula--> x= [-b +-( sqrt b^2-4ac)]/2a a=6 b= -5 c= -3 x= 5+- [sqrt(-5^2-4(6)(-3)/2(6) 5+-sqrt(25+72)/12 5+-sqrt(97/12)

OpenStudy (shamim):

U should not give the whole solution in 1 post @hockeychick23